OpenStudy (anonymous):

Find the center and radius of a circle whose equation is x^2+y^2+88=8y-22x??? Can someone please help me??? I have no idea how to do this!!

OpenStudy (aum):

Group the "x" terms together and complete the square. Group the "y" terms together and complete the square. Express result in the form: (x-h)^2 + (y-k)^2 = r^2 Center will be: (h,k) and radius will be r.

OpenStudy (anonymous):

if the equation is of the from x^2+y^2 +2gx +2fy +c =0 then centre of the circle is (-g,-f) and radius is sqrt(g^2+f^2 -c) compare the given x^2+y^2+88=8y-22x or ( x^2+y^2 +22x -8y +88=0) with the x^2+y^2 +2gx +2fy +c =0 and identify f,g,c

OpenStudy (anonymous):

I still don't understand... What am I supposed to do to get the answer?

OpenStudy (anonymous):

here u have 22 =2g means g =11 and 2f =-8 means f=-4 also c is 88 so centre is (-11,4) and radius is sqrt( 11^2 +4^2+88) =sqrt(225) = 15